# WarmServe: Enabling One-for-Many GPU Prewarming for Multi-LLM Serving

### v1

@@ 3 WarmServe Overview
WarmServe has no impact on inference performance since it allows serving instances to have exclusive access to GPUs. Meanwhile, it adopts the idea from GPU sharing systems [li2023alpaserve, muxserve, prism, qlm] to colocate prewarmed models on universal GPU workers. This approach maximizes prewarming efficiency and significantly reduces the TTFT for cold-start models since the model preparation process is launched in advance.
4 WarmServe Universal GPU Workers
WarmServe classifies GPU workers in the cluster into three categories: idle, universal, and dedicated. While dedicated GPU workers run inference tasks for specific models, universal workers store weights from multiple models simultaneously. This allows workers to quickly transition into serving instances for any of these models. To prewarm an LLM, its weights are loaded into several selected idle GPU workers within the same server. These GPUs are then able to immediately begin inference upon scaling up.
Figure 5(a) illustrates the lifecycle of a GPU worker. WarmServe loads weights from different LLMs into idle GPU workers to transform them into universal workers. When a model experiences a load spike, these workers can be exclusively assigned to that model after evicting all other models. At this point, the workers become dedicated GPU workers, with the remaining GPU memory allocated for the KV cache.
To reduce the size of loaded weights, we avoid prewarming the entire model across universal GPU workers. Instead, we prewarm only the first several layers and perform parallel model loading and inference when spawning the instance, similar to prior works [bai2020pipeswitch, paraserve, blitzscale, lambdascale]. Specifically, we conduct an offline profiling pass to obtain the number of layers that must be preloaded to ensure that inference is not blocked. These numbers are then directly applied during online prewarming.
@@ 4.1 Proactive Prewarming
WarmServe leverages proactive prewarming to quickly prewarm models under dynamic LLM workloads. As described in §2.1, instances enter a grace period before final termination. During this period, we proactively load weights from different models into the GPUs of these instances. While they continue processing ongoing requests, we utilize their unused KV cache space to store new parameters. This approach is viable because these instances are typically underutilized and no longer receive additional requests.
Figure 5(b) illustrates the proactive prewarming process. For dedicated GPU workers in grace periods, we monitor the amount of free KV cache space. If sufficient memory is available, we load the weights of other models into unused KV cache to prewarm new models. Once the instance is terminated, these workers transition directly into universal GPU workers that contain both the previously running model and the proactively loaded models. This quick conversion allows us to prepare universal GPU workers in advance, ensuring prewarming completion before the next allocation.
However, determining the appropriate amount of free KV cache space to allocate for prewarming new models is challenging. Since these dedicated GPU workers are still processing ongoing requests, their KV cache requirements can increase over time. For example, a single request with a sequence length of 128K can consume the majority of the KV cache space when serving a Llama3.1-8B model with a single A100-40G GPU [llama3]. Consequently, an aggressive prewarming strategy that provides too much memory for prewarming may lead to insufficient KV cache space, forcing the eviction of the recently loaded weights. Conversely, a conservative prewarming strategy can cause the underutilization of GPU memory, slowing down the conversion into universal workers.
Given that the total generation length of an LLM request is commonly regarded as unpredictable [fastserve, superserve], WarmServe strikes a balance between aggressive and conservative prewarming strategies. Formally, let represent the number of ongoing requests and be the maximum batch size of the instance. The GPU utilization at the request granularity can be computed as . Furthermore, let denote the total KV cache capacity and be the currently consumed portion. The amount of GPU memory to reserve for prewarming can be determined using the following formula.
Eq. 1 consists of two parts: (1) the expected KV cache usage based on the number of ongoing requests and (2) the actual KV cache usage plus a reserved buffer for potential KV space expansion. Since the total generation length is unpredictable, we allocate additional KV cache space for ongoing requests by estimating the average KV cache usage per request, calculated as .
Upon the completion of each request for a grace-period instance, WarmServe verifies whether the size of the current reserved KV cache exceeds the reservation target. If it does, the surplus KV blocks are released to prewarm new models. The instance then reports the amount of newly freed memory to the global manager via a shared memory region.
In addition to proactive prewarming, WarmServe ensures that even if prewarming has not completed upon instance startup, the partially prewarmed weights are immediately operational. In this case, the instance can directly begin inference on the available layers while asynchronously loading the remaining parameters in the background.
4.2 Zero-Overhead Memory Switching
@@
WarmServe uses a zero-overhead memory switching strategy to efficiently switch GPU memory among different LLMs. The strategy leverages the CUDA Virtual Memory Management (VMM) API [cuda-vmm] to directly manipulate the GPU page table, similar to previous works [prism, kunserve, vattention]. This API allows us to pre-allocate all physical pages in the GPU and map virtual memory addresses to them on demand.
@@
Our method begins by pre-allocating multiple virtual memory regions on each GPU, where the size of each region is equivalent to the total available GPU memory. Each of these regions serves as a prewarm slot that is used to store an individual prewarmed model. During the prewarming phase, we assign models different slots and allocate the necessary physical pages for each model according to model sizes. Subsequently, the corresponding virtual addresses within the slot are mapped to these physical pages, with model weights always placed at the start of its respective slot.

### v2

@@ 3 Method
WarmServe avoids degrading steady-state inference performance by ensuring exclusive GPU access for serving instances. Meanwhile, it adopts the idea from GPU sharing systems (Li et al., 2023; Duan et al., 2024; Yu et al., 2025c; Patke et al., 2024; Gao et al., 2025; Xiang et al., 2025) to colocate prewarmed models on universal GPU workers. This approach maximizes prewarming efficiency and significantly reduces the TTFT for cold-start models since the model preparation process is launched in advance.
3.1 Workload Prediction
Problem Formulation. We partition time into windows of length minutes. The workload predictor aims to forecast, for each model, the average and peak load in the upcoming time window given historical invocations.
Input. Taking peak load as an example, the historical invocations are denoted as , which represents the maximum concurrent requests of the LLM in the -th time window on day .
Output. The predicted peak load for the -th window on day , denoted as .
Prediction Algorithm. WarmServe predicts the workload for the next time window using a simple corrective seasonal predictor (CSP) (Holt, 2004; Winters, 1960; Ghosh et al., 2009). The prediction is based on two components: a seasonal pattern and a corrective pattern. For the seasonal pattern, we use the average peak load of the -th window in the past several days, calculated as follows.
where is the seasonal term for day , window . is the number of historical days to sample, typically set to .
For the corrective pattern, we utilize recently observed peak loads to quantify the deviation between the current trend and the historical seasonal pattern, calculated as follows.
where is the corrective term for day , window . is the size of the lookback window, typically set to . This weighting scheme gives more importance to recent errors.
The prediction of average load follows the same procedure. Due to the periodic nature of LLM workloads, our predictor demonstrates high performance on real-world datasets. With a 5-minute window size, it achieves an average relative error of 5.3% for average loads and 7.3% for peak loads on the AzureConv (Stojkovic et al., 2025) trace. While more sophisticated prediction algorithms, such as ARIMA (Box et al., 2015) and deep learning models (Ni et al., 2026; Liu et al., 2017), could potentially yield higher accuracy, we found that CSP is sufficiently effective for guiding model prewarming while imposing negligible overhead.
3.2 Model Placement
Input: #models , #GPUs ; model size , parallelism degree , batch size , current number of instances , predicted average and peak load in the next time window , cold start latency for each model ; available memory for each GPU .
Output: to-prewarm models and placement .
Problem Formulation. Based on prediction results, WarmServe prewarms models on GPUs that have available memory. During this process, we need to determine which model to prewarm and where to store its weights.
Input. Consider the cluster has GPUs, where the -th GPU has free memory. Assume we have models in total, where the -th model already has active serving instances. For each model, the predictor will give the predicted average load and peak load in the next time window. Each model also has characteristics including model size , required number of GPUs , and maximum batch size .
Output. A list of models to prewarm and the set of GPUs to place each model in. Note that the model list can be duplicated since a model can have multiple prewarmed replicas to prepare for burst requests.
Placement Algorithm. For the -th model, we want to prewarm at most replicas to handle future requests. WarmServe first computes a prewarming score for each replica that represents its necessity. The score is determined by (1) the gap between the number of active instances and the predicted future loads and (2) the time cost to launch an instance from scratch.
Based on prewarming scores, WarmServe sorts replicas by their scores in descending order, then iterates through replicas and attempts to place each one. For each replica, we first derive several GPU groups that can serve as candidate placement positions, and then greedily select the optimal one. The selection process tries to isolate high-score replicas to prevent mutual interference and place lower-score replicas in a way that minimizes their impact on existing replicas. Specifically, we prioritize groups where the new replica’s score is higher than that of any other existing replica nested within the group. If multiple such groups are available, the one with the minimum sum of scores from its nested replicas is chosen. Otherwise, the algorithm defaults to selecting the group with the minimum sum of scores. Algorithm 1 outlines the model placement algorithm. For a more detailed explanation, please refer to Appendix B.
@@ 3.3 Proactive Prewarming
WarmServe leverages proactive prewarming to quickly prewarm models under dynamic LLM workloads. Typically, an LLM serving cluster adopts an autoscaler to scale the number of serving instances for each model. If the GPU utilization of a model falls below a predefined threshold, the autoscaler attempts to shut down some of its instances. At this time, WarmServe proactively loads weights from different models into the GPUs of these to-be-terminated instances. While they continue processing existing requests, we utilize their unused KV cache space to store new parameters, as shown in Figure 3. These instances have plenty of unused KV cache space since they are usually underutilized and no longer receive additional requests.
Problem Formulation. Dedicated GPUs that will be released in the near future should inform the global manager how much KV cache space they can provide to store new models. As these GPUs are still processing ongoing requests, they must determine how much KV cache space to reserve for existing requests.
Input. Denote as the number of existing requests and as the maximum batch size of the instance. Assume the total KV cache capacity is and is the size of currently consumed space.
Output. The amount of KV cache space to be reserved for existing requests, denoted as .
Reservation Strategy. We consider the reservation problem with two targets: (1) the expected KV cache usage based on the number of ongoing requests and (2) the current KV cache usage plus a reserved buffer for potential KV space expansion. The reserved KV cache capacity is computed as the maximum of two targets:
Since generation lengths are unpredictable, we reserve extra KV cache for ongoing requests using the estimated average per-request usage . When the reserved KV cache space becomes insufficient to serve existing requests, we evict prewarmed weights to free up memory.
3.4 GPU Memory Management
@@
WarmServe uses a memory switching mechanism to efficiently manage GPU memory among different prewarmed LLMs. The mechanism leverages the CUDA Virtual Memory Management (VMM) API (NVIDIA, 2025b) to manipulate the GPU page table, similar to previous works (Yu et al., 2025c; Cheng et al., 2026; Prabhu et al., 2025). This API allows us to pre-allocate all physical pages in the GPU and map virtual memory addresses to them on demand.
@@
Our method begins by pre-allocating multiple virtual memory regions on each GPU, where the size of each region equals the total available GPU memory. Each of these regions serves as a prewarm slot that is used to store an individual prewarmed model. During the prewarming phase, we assign models different slots and map physical pages to them. Models can load their weights into corresponding slots. When a prewarmed model needs to occupy the GPU and become a serving instance, we map physical KV cache space to its prewarm slot and invalidate all other prewarm slots. The serving engine will use this prewarm slot to perform all computations and only access the serving model. In this way, a universal GPU can switch between different prewarmed models upon load spikes.
All page table modifications are overlapped with other operations so that the overhead of our GPU memory switching mechanism is negligible. More details are in Appendix C.
Implementation. WarmServe is implemented based on vLLM (Kwon et al., 2023). We use Ray Actor (Moritz et al., 2018) to instantiate GPU workers and perform an end-to-end prewarming during serving engine creation. The code of WarmServe is open-source and is publicly available at https://github.com/LLMServe/WarmServe. For more implementation details, please refer to Appendix D.
4 Evaluation
| Model | Size (GB) | # Required GPUs | Trace Sampling Day |
| Llama2-7B-0 | 12.55 | 1 | Thursday |
| Llama2-7B-1 | 12.55 | 1 | Friday |
| Llama2-13B | 24.24 | 2 | Saturday |
| Llama2-70B | 128.49 | 4 | Sunday |
| Model | TTFT (s) | ||
| No Prewarm. | With Prewarm. | Warm | |
| Llama2-7B | 25.30 | 0.32 (79.06) | 0.05 (506.0) |
| Llama2-13B | 28.04 | 0.32 (87.63) | 0.05 (560.8) |
| Llama2-70B | 40.29 | 0.67 (60.13) | 0.05 (805.8) |
In this section, we present experimental results to validate the efficiency and effectiveness of WarmServe. We also perform ablation studies to verify the effectiveness of individual components. For additional experimental results, please refer to Appendix A.
4.1 Experimental Setup
Testbed. We evaluate WarmServe on two GPU servers, each with eight GPUs. A single GPU has around 1K TFLOPS of computational power for dense FP16 computation and intra-server GPUs are connected via 400 GB/s NVLink. Each server is equipped with eight 200 Gbps RDMA NICs for inter-server connectivity. Model weights are stored in host memory and are loaded into GPUs on demand via PCIe 5.0 x16 channels.
Metrics. We primarily focus on the time-to-first-token (TTFT) in the experiments, defined as the end-to-end duration from request submission to receipt of the first token. We also provide experimental results regarding the time-per-output-token (TPOT) in the appendix.
Workloads. Following a similar approach to previous works (Fu et al., 2024; Duan et al., 2024; Lou et al., 2026; Xiang et al., 2025), we use the Llama2 model series (Meta, 2023) and duplicate Llama2-7B to increase the scaling frequency. The details of models are shown in Table 1. With respect to memory capacity, a single GPU can store the weights of six whole Llama2-7B models.
Following prior work (Zhang et al., 2025), we generate workloads from the AzureConv trace (Stojkovic et al., 2025), sampling model requests from different days in the trace. Per-model request rates follow a power-law distribution with exponent , as in (Duan et al., 2024). We vary the global requests-per-second (RPS) to control the total load.
Baselines. We compare WarmServe with the following baselines.
ServerlessLLM-GPU (SLLM-GPU) (Fu et al., 2024): Since model weights already exist in host memory, we extend the caching mechanism of ServerlessLLM to GPUs. When an instance stops, its model parameters remain in GPU memory for future invocation.
MuxServe (Duan et al., 2024): It colocates models on GPUs and uses CUDA MPS (NVIDIA, 2025a) to isolate inference tasks. LLM serving instances are created in advance. As the original MuxServe implementation is based on an old vLLM version, we reimplement it using the same vLLM version as WarmServe for fairness.
All systems use vLLM (Kwon et al., 2023) as the inference backend. SLLM-GPU and WarmServe adopt a batch size of 32, while MuxServe uses the configuration generated by itself in each setting. We do not compare with BlitzScale (Zhang et al., 2025) since its documentation is incomplete and we are unable to install it in our environment.
4.2 Prewarming Effectiveness
We first evaluate the effectiveness of prewarming by measuring the TTFT of models under different scenarios. Table 2 shows the TTFT of models under (1) No Prewarming, (2) With Prewarming, and (3) Warm. In the first two cases, the serving engine is created upon request arrival, and model weights are previously stored in host memory and GPU memory, respectively. In the Warm case, the engine is already running and can process requests immediately.
The results demonstrate that successful prewarming can significantly reduce TTFT. With prewarming, WarmServe can achieve a 60.13–87.63 TTFT reduction. In this case, the first token can be generated in 670ms for a Llama2-70B model when the serving engine is created on demand. This is sufficient to satisfy most service level objectives for LLM serving in production.
4.3 End-to-End Experiments
We further evaluate the effectiveness of WarmServe through end-to-end experiments. Figure 4 presents the tail TTFT of systems under various scenarios. To validate the effectiveness of proactive prewarming, we also conduct experiments on WarmServe with proactive prewarming disabled.
The results show that WarmServe consistently delivers low TTFT across all settings. Compared to SLLM-GPU, WarmServe achieves a 1.07–10.06 reduction in P95 TTFT and a 1.53–50.79 reduction in P99 TTFT by rapidly launching new instances from prewarmed models. Moreover, due to the dynamic nature of workloads, proactive prewarming drastically improves prewarming efficiency, reducing tail TTFT by 1.03–32.87. The tail TTFT in autoscaling-based systems remains relatively stable, as higher traffic only increases the frequency of scaling rather than the percentage of requests that experience delays.
The GPU-sharing system, MuxServe, exhibits performance comparable to WarmServe under light loads due to its pre-created model instances. However, its static model placement strategy limits serving capacity and degrades performance for colocated models, leading to severe queuing under heavy loads. For example, at an RPS of 15 and =0.5, MuxServe’s P95 and P99 TTFT are 34.5 and 134.0 higher than those of WarmServe, respectively. Overall, WarmServe can handle up to 2.5 more requests than MuxServe while maintaining low TTFT.
4.4 Ablation Study
We conduct an ablation study to analyze the impact of the prediction window size and validate the effectiveness of our key components: model prewarming, the model placement algorithm, and the proactive prewarming strategy. We set RPS to 25 and use a 5-minute window size by default.
Figure 5 shows that disabling model prewarming, the model placement algorithm, or the proactive prewarming strategy reduces the percentage of requests meeting the 100ms TTFT threshold to 0.15, 0.29, and 0.88 of the baseline, respectively. These results highlight that while the prewarming mechanism itself is fundamental to performance, our model placement algorithm and proactive prewarming strategy provide a critical additional boost.
The prediction window size involves a critical trade-off: small windows (e.g., 3 min) suffer from prediction instability, while overly large windows (e.g., 40 min) fail to capture transient load dynamics. These settings serve only 0.46 and 0.30 as many requests within the 100ms threshold compared to the 5-minute baseline. Notably, even with suboptimal windows, WarmServe consistently outperforms the no-prewarming baseline, demonstrating the robustness of our GPU prewarming mechanism.
| RPS | WarmServe | SLLM-GPU | MuxServe | Pred. Autoscal. |
| 320 | 0.18 | 5.1 (28.3) | 0.27 (1.5) | 0.18 (1.0) |
| 480 | 0.25 | 5.1 (20.4) | 35 (140) | 0.25 (1.0) |
| 640 | 0.40 | 12 (30.0) | 400 (1000) | 1.84 (4.6) |
| 800 | 0.38 | 8.9 (23.4) | 1300 (3400) | 4.42 (11.6) |
4.5 Large-Scale Simulation
We conduct large-scale simulations to further evaluate the scalability of WarmServe. The simulation is based on the same workload generation method as in the end-to-end experiments (), but with a larger cluster of 512 GPUs. To show the necessity of prewarming, we also include a baseline that uses the same workload predictor as WarmServe to perform predictive autoscaling without prewarming.
The simulation results are shown in Table 3. WarmServe consistently achieves low tail TTFT across all load levels, demonstrating its high scalability. Conversely, the performance of SLLM-GPU and MuxServe degrades significantly as the load increases, with TTFT increasing by up to 30 and 3400, respectively. Predictive autoscaling can achieve the same TTFT as WarmServe under light loads through timely autoscaling. However, as load increases, the frequency of scaling events rises, leading to more requests experiencing scaling delays. Owing to the one-for-many GPU prewarming mechanism, WarmServe can leverage limited spare resources to prepare for future demands and maintain low TTFT even under heavy loads.
4.6 Workload Prediction
We evaluate our workload predictor, CSP, by predicting workload characteristics for a single model using production traces including AzureConv, AzureCode, BurstGPT, and ServeGen. We utilize the first week of data from each trace, partitioned into 5-minute windows. For every window between Tuesday and Sunday, CSP predicts the average and peak loads based on the historical data of all preceding windows. Figure 6 shows the hourly prediction performance across these traces. Throughout the evaluation period, the predicted loads align closely with the ground truth, demonstrating the strong long-term predictability inherent in production LLM traces.
5 Related Work
Multi-LLM Serving. Existing systems generally follow two paradigms: autoscaling, which allocates exclusive GPUs to instances (Fu et al., 2024; Yu et al., 2025a; Hu et al., 2025; Zhang et al., 2025; Yu et al., 2025b; Zhu et al., 2025; Liu et al., 2025; Lou et al., 2026; Mei et al., 2026), and GPU sharing, which colocates multiple models to increase utilization (Li et al., 2023; Duan et al., 2024; Yu et al., 2025c; Patke et al., 2024; Gao et al., 2025; Xiang et al., 2025). P-LoRA (Ni et al., 2026) prefetches LoRA adapters. However, its scope is limited to lightweight LoRA modules rather than full-parameter models. Autoscaling approaches often optimize model loading via caching (Fu et al., 2024; Yu et al., 2025a; Hu et al., 2025; Zhu et al., 2025; Liu et al., 2025), high-bandwidth transfers (Zhang et al., 2025; Yu et al., 2025b), or distributed fetching (Lou et al., 2026). In contrast, WarmServe adopts a proactive autoscaling approach and prewarms models using predictions.
Serverless Prewarming. Prewarming has been widely used to reduce cold start latency in serverless computing (Bhasi et al., 2021; Gunasekaran et al., 2020; Stojkovic et al., 2023; Cai et al., 2024; Shahrad et al., 2020; Sahraei et al., 2023; Yu et al., 2024; Du et al., 2020). However, LLMs present unique challenges: they span multiple GPUs and require full weight loading. WarmServe tailors prewarming placement and memory management for LLM serving.
KV Cache Management. Efficient KV cache management is critical for LLM performance. While prior works optimize KV cache management via paged memory (Kwon et al., 2023), long-sequence handling (Lin et al., 2024; Wu et al., 2024, 2025), or offloading (Qin et al., 2025; Sheng et al., 2023), WarmServe repurposes unused KV cache memory to store prewarmed model weights, enabling a GPU worker to transition roles near-instantaneously.
6 Conclusion
This paper presents WarmServe, a multi-LLM serving system that leverages the long-term predictability of LLM workloads to enable one-for-many GPU prewarming. WarmServe employs a specialized model placement strategy to mitigate cross-model prewarming interference, and proactively loads models onto active GPUs to speed up prewarming. Evaluation results show that WarmServe drastically improves inference performance compared to existing systems.
Acknowledgments
We sincerely thank the anonymous reviewers for their valuable feedback on this paper. This work was supported in part by the National Key Research and Development Program of China under Grant 2022YFB4500700, the Scientific Research Innovation Capability Support Project for Young Faculty under Grant ZYGXQNJSKYCXNLZCXM-I1, and the National Natural Science Foundation of China under Grant 62172008 and 62325201. Xin Jin and Chen Sun are the corresponding authors. Chiheng Lou, Sheng Qi, Xuanzhe Liu, and Xin Jin are also with the Key Laboratory of High Confidence Software Technologies (Peking University), Ministry of Education.
Impact Statement
This paper presents work whose goal is to advance the field of Large Language Model Serving Systems. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Additional Experimental Results
A.1 TTFT for Different Models
Figure 7 details the TTFT for different models under RPS=25 in our experiments. For =0.5, the MuxServe-chosen placement colocates the 7B-0 and 70B models on 8 GPUs, whereas for , the 7B-0, 7B-1, and 13B models all share the same 8 GPUs. Although MuxServe can achieve low TTFT for models that do not share GPUs, colocated models are constrained by limited GPU resources, leading to severe request queuing. GPU sharing also introduces performance interference. For example, under the =2.0 setting, the 13B model receives only 8.2% of total requests but still experiences significant queuing. This is due to resource contention from the high-demand 7B-0 model, which processes 70.2% of the requests.
A.2 TPOT for Different Systems
We provide the TPOT CDF of systems in Figure 8 and Figure 9. MuxServe’s strategy of increasing model parallelism and colocating multiple models on the same GPUs leads to severe inference performance degradation. For example, under heavy load (RPS=25) in Figure 8, while over 60% of requests in autoscaling-based systems achieve a TPOT under 50ms, the overhead from GPU sharing causes over 40% of MuxServe’s requests to exceed a TPOT of 100ms. This analysis highlights the fundamental advantage of WarmServe: its prewarming techniques drastically reduce TTFT, and its exclusive allocation of GPU workers preserves inference performance.
A.3 Evaluation on AzureCode Dataset
Figure 11 and Figure 11 depict the TTFT and TPOT of systems on the AzureCode trace, with fixed at 0.5. This trace contains fewer requests than AzureConv, making it less predictable. Consequently, WarmServe exhibits a slight performance degradation on this trace.
A.4 Prewarming Hit Ratio
Figure 12 shows the effectiveness of one-for-many prewarming by presenting the prewarming hit ratios of WarmServe. Under light loads, the cluster has sufficient idle GPUs, allowing WarmServe to prewarm all required model replicas and ensure a 100% prewarming hit ratio. As the load increases, the resources available for prewarming are reduced. Despite this, WarmServe successfully prewarms the majority of required instances, achieving an average hit ratio of 82% under RPS=25. This high hit ratio confirms the viability of one-for-many prewarming under pressure.
Appendix B Model Placement Algorithm
B.1 Prewarming Score Computation
To compute the score of each prewarming replica, we classify replicas for each model into two categories.
Burst replicas. These replicas are used to tackle load spikes. After creating sufficient basic replicas, we continue to prewarm additional replicas until the total number of instances (both active and prewarmed) can serve the predicted peak load. These additional replicas are designated as burst replicas, ensuring enough serving capacity under peak loads.
Formally, consider a model with active instances and a batch size of , and with predicted average and peak loads and , respectively. The number of basic () and burst () replicas to be prewarmed is calculated as follows.
For burst replicas, the score (Eq. 8) contains an additional burstiness factor, . This term represents the burstiness of the load in the upcoming time window, ensuring that models expecting a larger spike in traffic are prioritized accordingly. Note that we make sure basic replicas always have higher priorities than burst replicas, regardless of their scores.
B.2 Placement Algorithm
The placement algorithm of WarmServe is governed by two primary guidelines. The first guideline strictly prohibits partial GPU sharing. Specifically, for any two models, the set of GPUs allocated to them must either be entirely disjoint or one set must be a complete subset of the other. This constraint ensures that prewarming contention for GPU resources is limited to models in a nested arrangement, thereby reducing the interference patterns that arise from more complex overlaps.
After calculating prewarming scores, our placement algorithm strategically allocates prewarming replicas to GPU workers. It begins by calculating the prewarming score for each replica and then proceeds to place them according to specific placement guidelines. Replicas are processed in descending order of their prewarming scores, with basic replicas prioritized over burst replicas. For each replica , the following procedure is executed.
Placement group formation. Next, from the pool of candidate GPU workers, the algorithm attempts to form valid placement groups. A valid group must meet two requirements: (1) all workers in the group must be located on the same server to guarantee inference performance, and (2) the group’s workers must not partially overlap with any other existing prewarming replica. If no valid group can be formed, we move on to the next replica.
Optimal group selection. If one or more valid groups exist, the algorithm greedily selects the optimal one. The selection process prioritizes groups where the new replica’s score is higher than any other existing replica that is nested within the group. If multiple such groups are available, the one with the minimum sum of scores from its nested replicas is chosen. Otherwise, the algorithm defaults to selecting the group with the minimum sum of scores.
Appendix C GPU Memory Switching Mechanism